# Pre-DPO: Improving Data Utilization in Direct Preference Optimization Using a Guiding Reference Model

### v2

@@ B.1 Response length differences
Normalized Length $\text{Normalized Length Difference}=\frac{|\text{len}(y^{+})-\text{len}(y^{-})|}{\max\left(\text{len}(y^{+}),\text{len}(y^{-})\right)},$ (11)
@@
We compute this metric on three datasets: the original UltraFeedback preference data used for Base models, and the resampled UltraFeedback-Llama3.2 and UltraFeedback-Qwen2.5 preference datasets used for Instruct models. The original UltraFeedback dataset exhibits an average normalized length difference of 0.465, while the resampled on-policy datasets, UltraFeedback-Llama3.2 and UltraFeedback-Qwen2.5, have smaller values of 0.296 and 0.224, respectively. This indicates that the resampled datasets contain less variation in response length, potentially imposing more stringent optimization conditions for SimPO.
B.2 The $\lambda$ distribution after training of Pre-DPO and vanilla DPO / SimPO
(a) Qwen2.5-7B-Base
(b) Qwen2.5-7B-Instruct
(c) Llama3.2-3B-Base
(d) Llama3.2-3B-Instruct
(e) Qwen2.5-7B-Base
(f) Qwen2.5-7B-Instruct
Figure 4: More violin plots showing the distribution of $\lambda$ values on the training preference dataset for Pre-DPO and DPO / SimPO across different settings (with $\beta=0.01$).
More differences in $\lambda$ distributions on the training preference dataset between Pre-DPO and DPO / SimPO are shown in Figure 4. Similar to the case where DPO is used as the guiding reference model, Pre-DPO also produces higher mean and median $\lambda$ values when using SimPO as the guiding reference.

### v3

@@ C.1 Response Length Differences of $y^{+/-}$ Across Datasets
$\text{Normalized Length Difference}=\frac{|\text{len}(y^{+})-\text{len}(y^{-})|}{\max\left(\text{len}(y^{+}),\text{len}(y^{-})\right)},$ (16)
@@
We compute this metric on three datasets: the original UltraFeedback preference data used for Base models, and the constructed UltraFeedback-Llama3.2-3B-Instruct and UltraFeedback-Qwen2.5-7B-Instruct preference datasets used for Instruct models. The original UltraFeedback dataset exhibits an average normalized length difference of 0.465, while the resampled on-policy datasets, UltraFeedback-Llama3.2-3B-Instruct and UltraFeedback-Qwen2.5-7B-Instruct, have smaller values of 0.296 and 0.224, respectively. This indicates that the resampled datasets contain less variation in response length, potentially imposing more stringent optimization conditions for SimPO.
C.2 One Iteration of Pre-DPO Can Be Enough
Method Initial Policy Ref. AlpacaEval 2
LC (%) WR (%) Len.
DPO SFT SFT 10.5 12.0 2042
Pre-DPO round 1 SFT DPO 12.5 13.9 2061
Pre-DPO round 2 SFT Pre-DPO round 1 12.2 14.0 2134
Table 5: A further round of Pre-DPO under the Llama3.2-3B-Base setting. With sufficient hyperparameter tuning, one round of Pre-DPO can be enough.
The purpose of the guiding reference model—whether derived from DPO or SimPO—is not to directly achieve optimal performance, but to facilitate more effective use of existing preference data and guide subsequent preference optimization. Additional empirical results indicate that, with proper hyperparameter tuning, a single iteration of Pre-DPO is generally sufficient for effective data utilization. As shown in Table 5, applying Pre-DPO for an additional iteration yields negligible improvements on AlpacaEval 2 in the Llama3.2-3B-Base setting.